# PuzzleWorld: A Benchmark for Multimodal, Open-Ended Reasoning in Puzzlehunts

### v1

@@ Abstract
Puzzlehunts are a genre of complex, multi-step puzzles lacking well-defined problem definitions. In contrast to conventional reasoning benchmarks consisting of tasks with clear instructions, puzzlehunts require models to discover the underlying problem structure from multimodal evidence and iterative reasoning, mirroring real-world domains such as scientific discovery, exploratory data analysis, or investigative problem-solving. Despite recent progress in foundation models, their performance on such open-ended settings remains largely untested. In this paper, we introduce PuzzleWorld, a large-scale benchmark of 667 puzzlehunt-style problems designed to assess step-by-step, open-ended, and creative multimodal reasoning. Each puzzle is annotated with the final solution, detailed reasoning traces, and cognitive skill labels, enabling holistic benchmarking and fine-grained diagnostic analysis. Most state-of-the-art models achieve only 1-2% final answer accuracy, with the best model solving only 14% of puzzles and reaching 40% stepwise accuracy. To demonstrate the value of our reasoning annotations, we show that fine-tuning a small model on reasoning traces improves stepwise reasoning from 4% to 11%, while training on final answers alone degrades performance to near zero. Our error analysis reveals that current models exhibit myopic reasoning, are bottlenecked by the limitations of language-based inference, and lack sketching capabilities crucial for visual and spatial reasoning. We release PuzzleWorld at https://github.com/MIT-MI/PuzzleWorld to support future work on building more general, open-ended, and creative reasoning systems.
@@ 1 Introduction
Figure 1: Overview of PuzzleWorld: PuzzleWorld is a dataset of complex puzzles that lack explicit instructions, requiring solvers to deduce the final answer from nuanced, multimodal cues from the puzzle content as well as external domain-specific knowledge. The raw puzzles and solutions are sourced from PuzzledPint, and the solutions, which are PNG images, are transcribed into a sequence of reasoning steps by human annotators. These annotations enable us to measure the accuracy of the final answer and the step-by-step progress made towards the solution. Best-viewed zoomed in and in color, high-resolution puzzles are in Appendix B.
Puzzles are designed to challenge a solver’s reasoning ability, creativity, and problem-solving skills. While some puzzles follow rigid formats—such as crosswords, Sudoku, or logic grids—others, like puzzlehunts, are intentionally open-ended. In puzzlehunts, solvers must first infer the nature of the task from ambiguous clues embedded in textual, visual, or cultural elements before identifying and systematically executing a multi-step solution path. These puzzles span diverse modalities and topics, including wordplay, logic, programming, visual reasoning, music, and pop culture. Often taking hours or days to solve, puzzlehunts are a staple at universities, hiring events, and community competitions, and have grown into a global phenomenon through both in-person and virtual formats.
@@
Beyond their entertainment value, puzzlehunts provide a high-ceiling challenge for evaluating human—and increasingly, artificial—reasoning. They demand compositional thinking, lateral reasoning, trial-and-error, and backtracking dead ends. Unlike most AI benchmarks that present well-specified tasks with clear instructions, puzzlehunts require solvers to discover both what the problem is and how to solve it. This makes them uniquely well-suited for evaluating general-purpose reasoning systems under conditions that more closely resemble real-world open-ended scenarios, such as scientific discovery, investigative analysis, or exploratory problem-solving.
Recent advances in language and multimodal reasoning [22] have enabled significant progress in step-by-step problem-solving [40, 44], transparent reasoning [10, 24], and enhanced human-AI collaboration [42, 6]. However, existing reasoning benchmarks largely focus on narrow domains with fully specified tasks, such as math [23] or coding [18]. These settings do not test a model’s ability to form hypotheses, adapt to implicit structures, or reason creatively across modalities. We argue that progress toward generalist AI requires new benchmarks that challenge models to operate in less structured, discovery-driven environments that demand flexible and holistic reasoning [26].
@@
To this end, we introduce PuzzleWorld, a benchmark of 667 real-world puzzlehunt problems curated from Puzzled Pint [29], a monthly puzzlehunt event series with content released under a Creative Commons license. These puzzles offer an open-ended, compositional challenge beyond prior benchmarks focused on instruction-following or task completion, and will grow with new puzzle releases over time. We collected fine-grained annotations of each puzzle for its final answer, the input modalities involved, the cognitive reasoning skills it exercises, and a manually curated step-by-step solution trace. These rich annotations support diagnostic analysis, model training, and detailed evaluation of intermediate reasoning capabilities.
@@
PuzzleWorld enables us to systematically study the multimodal and multi-step reasoning capabilities of today’s best foundation models. Most state-of-the-art models achieve only 1-2% final answer accuracy, with the best model solving only 14% of puzzles and reaching 40% stepwise accuracy. We additionally find that detailed annotations are important, as fine-tuning a model on annotated reasoning traces doubles a small model’s performance, while learning directly from final solutions can hurt the model and drive the accuracy towards zero. We also conduct detailed error analysis on models’ performance on PuzzleWorld, yielding tangible directions for future work in improving multimodal open-ended reasoning in AI. Together, these elements position PuzzleWorld as a rigorous resource for evaluating and improving general-purpose multimodal reasoning in AI systems. PuzzleWorld is publicly available at https://github.com/MIT-MI/PuzzleWorld. In the long run, we believe PuzzleWorld can catalyze more general and adaptable AI for mathematical and logical reasoning, open-ended scientific discovery, and autonomous assistive agents.
@@ 2 Related work
LLMs have demonstrated remarkable emergent capabilities, often matching or even surpassing human performance across a wide range of tasks [32]. Notably, models such as GPT-4 [2] and Claude [1] have achieved strong results not only on traditional NLP benchmarks—like question answering, summarization, and translation [41, 31, 30], but also in more complex domains such as mathematical reasoning, programming, and logical deduction [3, 18, 19]. These abilities suggest that LLMs are beginning to exhibit general-purpose reasoning skills, making them increasingly relevant to both academic research and practical applications.
This growing interest in reasoning evaluation has coincided with the rapid expansion of the LLM ecosystem, which includes both open and closed-source models. Open-source models such as LLaMA [36], Mistral [17], and Qwen [4] have become increasingly capable and accessible, while specialized variants like DeepSeek-R1 [14] are explicitly designed to improve reasoning by integrating reinforcement learning and other post-training strategies. Meanwhile, closed-source models such as GPT-o3 [28], Claude 3 Sonnet [1], and Gemini Thinking [12] have pushed the boundaries of real-world performance. However, despite these impressive capabilities, understanding the full extent and limitations of LLM reasoning remains a crucial open question, underscoring the need for benchmarks that rigorously assess their capability for flexible, holistic reasoning [26, 5].
@@
Numerous reasoning benchmarks have been proposed to evaluate various cognitive skills, including visual mathematical reasoning [23], spatial understanding [38], analogical reasoning [46], and social reasoning [20, 25]. However, few have addressed abstract, open-ended problems that demand holistic reasoning. HEMM [21], SciBench [39], MMMU [47], MMMU-Pro [48], MMT-Bench [45], and OlympiadBench [15] test multimodal reasoning across various disciplines in academic and real-world contexts. While these tasks are broad and challenging, they typically involve well-defined questions and are grounded in familiar domains that closely resemble the training distributions of large models. As such, they primarily assess in-distribution reasoning rather than creativity or adaptability. ARC-AGI [9] tests the ability to reason and adapt to new situations through abstract visual pattern recognition tasks that require minimal prior knowledge, yet it lacks the open-ended, exploratory nature of real-world problem solving. In contrast, PuzzleWorld  targets open-ended reasoning through puzzlehunts that lack explicit instructions or solution paths. Solving these tasks requires creatively piecing together subtle hints, often across many modalities, into coherent multi-step reasoning chains.
Figure 2: Overview of samples from PuzzleWorld Left: To gain a deeper understanding of model performance on PuzzleWorld, each puzzle is annotated with the input modalities of the puzzle content, the reasoning skills required to solve the puzzle, and step-by-step reasoning steps. Right: Example modality and reasoning skill annotations on three puzzles. High-resolution puzzle images are in Appendix B.
@@
A growing line of work has explored the use of puzzles to test the reasoning capabilities of AI systems. PuzzleVQA [8] consists of 2k puzzles that require abstracting patterns based on colors, numbers, shapes, and other concepts from visual puzzles to answer multiple-choice questions. AlgoVQA [13] is a benchmark of visual puzzles that require algorithmic reasoning. PUZZLES [11] tests the ability of RL agents to perform algorithmic reasoning on a set of 40 puzzles. While valuable for evaluating specific skills, these benchmarks focus on narrow domains with constrained task formats, and modern models generally perform quite well on these benchmarks [8, 27, 47]. On the other hand, the unstructured nature of the puzzlehunt problems in PuzzleWorld requires models to interpret ambiguous cues, explore creative solution strategies, and integrate information across diverse modalities and knowledge areas. These tasks often demand lateral thinking, symbolic abstraction, and visual-spatial reasoning - capabilities that go well beyond conventional pattern recognition or formal logic. The closest to our benchmark is EnigmaEval [37], which also evaluates AI’s reasoning capabilities on puzzles. However, EnigmaEval is a closed-source evaluation-only dataset and does not include manually annotated step-by-step solutions. The open-source puzzles and rich annotations in PuzzleWorld support fine-grained analysis of intermediate reasoning processes and common failure modes, facilitating the development and evaluation of more robust, general-purpose reasoning models.
@@ 3 Taxonomizing Multimodal Reasoning in Puzzlehunts
To understand how solving puzzlehunts engages reasoning capabilities evaluated separately in benchmarks like MMMU [48] and MathVista [23], we analyze puzzle solutions and classify them along two orthogonal dimensions: input modality and reasoning mechanism. This dual-axis taxonomy provides a comprehensive evaluation framework that captures both the form in which information is presented and the cognitive strategies required for reasoning.
@@ 3.1 Puzzle input modalities
To characterize the diverse nature of puzzles in our dataset, we categorize input content into three modalities: Text, which encompasses purely textual information such as written instructions, narratives, clues, or word puzzles, testing the model’s ability to comprehend and extract relevant linguistic information; Visual, which includes unstructured visual elements like images, icons, diagrams, and typography, challenging the model’s capacity to interpret visual semantics and patterns; and Structured, which refers to systematically organized visual information—such as tables, graphs, grids, matrices, and charts—requiring the model to understand relationships between data elements within defined structures. Table 1 shows the distribution of puzzles across modality and difficulty.
@@ 3.2 Puzzle reasoning mechanisms
We identify six core cognitive abilities essential for effective puzzle-solving in PuzzleWorld. These include logic, which covers deductive and inferential reasoning such as pattern recognition and causal inference; wordplay, involving flexible linguistic interpretation through puns, anagrams, and homophones; spatial reasoning, which tests an AI’s ability to mentally manipulate objects and navigate geometric structures; and cryptic decoding, which requires recognizing and applying transformations like ciphers and hidden encodings. In addition, knowledge-based reasoning leverages domain-specific facts from areas such as science or history, while commonsense reasoning draws on implicit real-world expectations.
This taxonomic approach enables quantitative analysis of puzzle complexity and facilitates targeted evaluation of AI reasoning capabilities across different cognitive dimensions. By mapping specific puzzles and reasoning tasks to combinations of modalities and mechanisms, we can identify areas of strength and weakness in AI systems, track progress over time, and guide future development efforts toward more balanced reasoning capabilities.
4 Creating PuzzleWorld
@@
Figure 3: Dataset construction procedure and statistics: Left: First, we source raw puzzles and solutions from PuzzledPint. As the PuzzledPint solutions are in PDF format and are often not correctly parsed by OCR (for example, some solutions consist of annotated figures rather than a textual description), the metadata and reasoning steps for each puzzle are human-annotated. We use GPT-4 to automatically flag puzzles that have unclear reasoning steps and incorrect flavor texts. Finally, two human verifiers perform a manual data cleaning on the flagged puzzles to ensure a consistent annotation format. Right: We summarize the statistics of our dataset. In particular, the average number of reasoning steps is high, and the steps are relatively complex, as shown by the high average word count.
@@ 4.1 Data collection and pre-processing
We collected our puzzle corpus from Puzzled Pint [29], an organization that publishes puzzles under the Creative Commons license (CC BY-NC-SA Intl. 4.0). Their repository contains monthly puzzle sets designed for collaborative solving in public venues, covering a diverse range of puzzle types and difficulties. This allowed us to obtain more than 700 raw puzzles spanning from 2010 to 2025.
Each puzzle in our dataset consists of its original PDF containing the puzzle content, a canonical single-phrase answer, and a corresponding solution document (either a text description or an annotated PDF explaining the solution process). Unlike Wang et al. [37], we deliberately preserved the original puzzle format rather than transcribing content into separate text and images. This decision was motivated by the observation that puzzle layouts often contain crucial spatial relationships and design elements integral to the solving process. Furthermore, Wang et al. [37] demonstrated that the best foundation models are not primarily constrained by OCR capabilities. Instead, we devote our manual annotation to construct fine-grained annotations of reasoning steps extracted from existing solution documents, ensuring that the annotations accurately capture the intended solution pathways while maintaining the integrity of the original puzzle presentation.

### v2

@@ Abstract
Puzzlehunts are a genre of complex, multi-step puzzles lacking well-defined problem definitions. In contrast to conventional reasoning benchmarks consisting of tasks with clear instructions and constrained environments, puzzlehunts requires discovering the underlying problem structure from multimodal evidence and iterative reasoning, mirroring real-world domains such as scientific discovery, exploratory data analysis, or investigative problem-solving. Despite progress in foundation models, their performance on open-ended settings remains largely untested. We introduce PuzzleWorld, a comprehensive benchmark of 667 puzzlehunt-style problems designed to assess step-by-step, open-ended, and creative multimodal reasoning. Each puzzle is annotated with the final solution, detailed reasoning traces, and cognitive skill labels, enabling holistic benchmarking and fine-grained diagnostic analysis. Most state-of-the-art models achieve only 1-4% final answer accuracy. On PuzzleWorld, the best model solves only 18% of puzzles and reaches 40% stepwise accuracy, matching human puzzle novices but falling significantly behind puzzle enthusiasts. To demonstrate the value of our reasoning annotations, we show that fine-tuning a small model on reasoning traces boosts stepwise accuracy from 4% to 11%, which translates to improvements in downstream visual reasoning tasks. Our detailed error analysis reveals that current models exhibit myopic reasoning, are bottlenecked by the limitations of language-based inference, and lack sketching capabilities crucial for visual and spatial reasoning. We release PuzzleWorld at https://github.com/MIT-MI/PuzzleWorld to support future work on building more general, open-ended, and creative reasoning systems.
footnotetext: *†These authors contributed equally.
@@ 1 Introduction
Recent advances in language and multimodal reasoning (Liang et al., 2024b) have enabled significant progress in step-by-step problem-solving (Wei et al., 2022; Yao et al., 2023), transparent reasoning (Creswell and Shanahan, 2022; Luo et al., 2023), and enhanced human-AI collaboration (Wu et al., 2022; Chen et al., 2025b). Such progress has been fuelled by and evaluated on comprehensive benchmarks, particularly in domains like mathematics (Lu et al., 2024) and code (Jiang et al., 2024). However, these benchmarks are largely confined to narrow, well-defined environments. In coding, tasks are meticulously specified and validated within executable environments (Jimenez et al., 2024). In geometry, models often rely on domain-specific languages to structure their reasoning (Chervonyi et al., 2025). While valuable, these benchmarks primarily test a model’s ability within a pre-defined problem space, rather than its ability to discover the problem itself.
In contrast, human reasoning excels in open-ended environments, where the rules are unstated and the objectives are ambiguous. We dynamically form hypotheses, adapt to implicit structures, and reason creatively across modalities to solve problems ranging from deciphering an escape room to novel scientific discovery. To build more generalist AI, we argue that the next frontier for evaluation lies beyond the current constrained settings. It demands benchmarks that challenge models to operate in less structured, discovery-driven environments that require more flexible and holistic reasoning (Mondorf and Plank, 2024).
Puzzles are designed precisely to test these human-like generalist reasoning abilities. While some popular puzzles such as Sudokus and crosswords are rigidly formatted, others, like puzzlehunts, are intentionally open-ended. In a puzzlehunt, puzzle solvers are not given a clear task; they must first infer the nature of the problem from ambiguous clues embedded in text, images, or cultural references before devising and executing a solution.
@@
Beyond their entertainment value, puzzlehunts model the essential challenges of real-world discovery and analysis. They demand compositional thinking, lateral reasoning, and the resilience to pursue leads, backtrack from dead ends, and manage uncertainty. Unlike current AI benchmarks that present well-specified tasks, puzzlehunts compel solvers to discover both what the problem is and how to solve it. This dual challenge makes them uniquely suited for evaluating general-purpose reasoning systems under conditions that more closely resemble open-ended scenarios like scientific investigation, intelligence analysis, or exploratory design.
@@
To bridge this gap, we introduce PuzzleWorld, a benchmark of 667 real-world puzzlehunt problems curated from Puzzled Pint (Puzzled Pint, 2025), a monthly puzzlehunt event with content released under a Creative Commons license. These puzzles offer an open-ended, compositional challenge beyond prior benchmarks focused on instruction-following or task completion, and will grow with new puzzle releases. For each puzzle, we provide fine-grained annotation of its solution, input modalities, cognitive reasoning skills it exercises, and a manually curated step-by-step solution trace. These rich annotations support diagnostic analysis, model training, and detailed evaluation of models’ reasoning capabilities. An overview of PuzzleWorld is provided in Figure 1.
@@
PuzzleWorld enables us to systematically study the multimodal and multi-step reasoning capabilities of today’s best foundation models. Most state-of-the-art models achieve only 1-4% final answer accuracy, with the best model solving only 14% of puzzles and reaching 40% stepwise accuracy. We additionally find that detailed annotations are important, as fine-tuning a model on annotated reasoning traces significantly improves a small model’s performance, both within PuzzleWorld and on other visual reasoning datasets. We also conduct detailed error analysis on models’ performance on PuzzleWorld, yielding tangible directions for future work in improving multimodal open-ended reasoning in AI. Together, these elements position PuzzleWorld as a rigorous resource for evaluating and improving general-purpose multimodal reasoning in AI systems. In the long run, we believe PuzzleWorld can catalyze more general and adaptable AI for mathematical and logical reasoning, open-ended scientific discovery, and assistive agents. PuzzleWorld is publicly available at https://github.com/MIT-MI/PuzzleWorld.
Figure 1: Overview of PuzzleWorld: PuzzleWorld is a dataset of complex puzzles that lack explicit instructions, requiring solvers to deduce the final answer from nuanced, multimodal cues from the puzzle content as well as external domain-specific knowledge. The raw puzzles and solutions are sourced from Puzzled Pint, and the solutions, which are PNG images, are transcribed into a sequence of reasoning steps by human annotators. These annotations enable us to measure the accuracy of the final answer and the step-by-step progress made towards the solution. Best-viewed zoomed in and in color, high-resolution puzzles are in Appendix E.3.
@@ 2 Related work
LLMs have demonstrated remarkable emergent capabilities, often matching or even surpassing human performance across a wide range of tasks (Street et al., 2024). Notably, models such as GPT-4 (Achiam et al., 2023) and Claude (Anthropic, 2025) have achieved strong results not only on traditional NLP benchmarks—like question answering, summarization, and translation (Widyassari et al., 2022; Soares and Parreiras, 2020; Singh et al., 2017), but also in more complex domains such as mathematical reasoning, programming, and logical deduction (Ahn et al., 2024; Jiang et al., 2024; Lam et al., 2024). These abilities suggest that LLMs are beginning to exhibit general-purpose reasoning skills, making them increasingly relevant to both academic research and practical applications. However, despite these impressive capabilities, understanding the full extent and limitations of LLM reasoning remains a crucial open question, underscoring the need for benchmarks that rigorously assess their capability for flexible, holistic reasoning (Mondorf and Plank, 2024; Chang et al., 2024).
@@
Numerous reasoning benchmarks have been proposed to evaluate various cognitive skills, including visual mathematical reasoning (Lu et al., 2024), spatial understanding (Wang et al., 2024a), analogical reasoning (Yiu et al., 2024), and social reasoning (Li et al., 2025; Mathur et al., 2025). However, few have addressed abstract, open-ended problems that demand holistic reasoning. HEMM (Liang et al., 2024a), SciBench (Wang et al., 2024b), MMMU (Yue et al., 2024b), MMMU-Pro (Yue et al., 2024a), MMT-Bench (Ying et al., 2024), and OlympiadBench (He et al., 2024) test multimodal reasoning across various disciplines in academic and real-world contexts. While these tasks are broad and challenging, they typically involve well-defined questions that closely resemble the training distributions of large models. As such, they primarily assess in-distribution reasoning rather than creativity or adaptability. ARC-AGI (Chollet, 2019) tests the ability to reason and adapt to new situations through abstract visual pattern recognition tasks that require minimal prior knowledge, yet it lacks the open-ended, exploratory nature of real-world problem solving. In contrast, PuzzleWorld targets open-ended reasoning through puzzlehunts that lack explicit instructions. Solving these tasks requires creatively piecing together subtle hints, often across many modalities, into coherent multi-step reasoning chains.
Figure 2: Overview of samples from PuzzleWorld. Left: To gain a deeper understanding of model performance on PuzzleWorld, each puzzle is annotated with the input modalities of the puzzle content, the reasoning skills required to solve the puzzle, and step-by-step reasoning steps. Right: Example modality and reasoning skill annotations on three puzzles. High-resolution puzzle images are included in Appendix E.3.
@@
A growing line of work has explored the use of puzzles to test the reasoning capabilities of AI systems. PuzzleVQA (Chia et al., 2024) consists of 2,000 puzzles that require abstracting patterns from visual puzzles to answer multiple-choice questions. AlgoVQA (Ghosal et al., 2024) is a visual puzzle benchmark requiring algorithmic reasoning. PUZZLES (Estermann et al., 2024) tests the ability of RL agents to perform algorithmic reasoning on a set of 40 puzzles. While valuable for evaluating specific skills, these benchmarks focus on narrow domains with constrained task formats, and modern models generally perform well on these benchmarks (Chia et al., 2024; Moskvichev et al., 2023; Yue et al., 2024b). On the other hand, the unstructured nature of the puzzlehunt problems in PuzzleWorld requires models to interpret ambiguous cues, explore creative strategies, and integrate information across diverse modalities and knowledge areas. In contrast to previous benchmarks that isolate strictly structured vertical reasoning (Chen et al., 2025a) or narrative-based lateral thinking (Huang et al., 2024), puzzlehunt puzzles require an integrated blend of lateral thinking, symbolic abstraction, and visual–spatial reasoning. The closest to our benchmark is EnigmaEval (Wang et al., 2025), which also evaluates AI’s reasoning capabilities on puzzlehunts. However, EnigmaEval is a closed-source evaluation-only dataset and does not include manually annotated step-by-step solutions. The open-sourced puzzles and rich annotations in PuzzleWorld support fine-grained analysis of intermediate reasoning and failure modes, facilitating the development and evaluation of more robust, general-purpose reasoning models.
@@ 3 Taxonomizing Multimodal Reasoning in Puzzlehunts
To understand how solving puzzlehunts engages reasoning capabilities evaluated separately in benchmarks like MMMU (Yue et al., 2024a), we analyze puzzle solutions and classify them along two dimensions: input modality and reasoning mechanism. This taxonomy provides a comprehensive evaluation framework that captures both the form in which information is presented and the cognitive strategies required for reasoning.
@@ 3.1 Puzzle input modalities
We consider three puzzle input modalities: Text, encompassing textual information such as instructions, narratives, or word puzzles, testing the model’s ability to extract relevant linguistic information; Visual, which includes unstructured visuals like images, icons, and typography, challenging the models to interpret visual semantics and patterns; and Structured, which refers to systematically organized visual information, such as tables, graphs, grids, matrices, and charts. Table 1 shows the distribution of puzzles across modality and difficulty.
@@ 3.2 Puzzle reasoning mechanisms
We identify six core cognitive abilities essential for effective puzzle-solving in PuzzleWorld. These include logic, which covers inferential reasoning such as deduction and causal inference; wordplay, involving flexible linguistic interpretation through puns, anagrams, and homophones; spatial reasoning, which tests an AI’s ability to mentally manipulate objects and navigate structures; and cryptic decoding, which requires recognizing and applying transformations like ciphers and hidden encodings. In addition, knowledge-based reasoning leverages domain-specific facts from areas such as science or history, while commonsense reasoning draws on implicit real-world expectations. This taxonomic approach enables targeted evaluation and analysis of AI reasoning capabilities across different cognitive dimensions. By mapping specific puzzles and reasoning tasks to combinations of modalities and mechanisms, we can identify areas of strength and weakness in AI systems, track progress over time, and guide future development efforts toward more balanced reasoning capabilities.
@@
Figure 3: Dataset construction procedure and statistics: Left: First, we source raw puzzles and solutions from Puzzled Pint. As the Puzzled Pint solutions are often not correctly parsed by OCR, each puzzle’s metadata and reasoning steps are human-annotated. We use GPT-4o to automatically flag ambiguous and inconsistent annotations. Finally, two human verifiers perform a manual data cleaning on the flagged puzzles to ensure a consistent annotation format. Right: We summarize the statistics of our dataset. The average number of reasoning steps is high, and the steps are relatively complex, as shown by the high average word count.
4 Creating PuzzleWorld
@@ 4.1 Data collection and pre-processing
We collected our puzzle corpus from Puzzled Pint (2025), an organization that publishes puzzles under Creative Commons (CC BY-NC-SA Intl. 4.0). Their repository contains monthly puzzles designed for collaborative solving, covering a diverse range of puzzle types and difficulties. This allowed us to obtain more than 700 raw puzzles spanning from 2010 to 2025.
Each puzzle in our dataset consists of its original PDF containing the puzzle content, a single-phrase answer, and a solution document. Unlike Wang et al. (2025), we deliberately preserved the original puzzle format rather than transcribing content into separate text and images. This decision was motivated by the importance of spatial relationships in puzzle layouts to the solving process. Furthermore, Wang et al. (2025) showed that the best foundation models are not primarily constrained by OCR capabilities. Instead, we devote our manual effort to construct fine-grained annotations of puzzle reasoning steps, ensuring that the annotations accurately capture the intended solution pathways while maintaining the integrity of the original puzzle presentation.